Adam is 5 times as old as charles, in 8 years time the sum of their ages will be equal to twice adam's present age. Find their present ages
Answer by mananth(16949)

Charles age =x
Adam age = 5x
...
8 years later
Charles age = x+8
Adam age = 5x+8
x+8+5x+8 = 2*5x
6x+16=10x
4x=16
x=4 Charles' age
Adam's age = 5x = 20 years